Electrons are accelerated by a potential difference of 100v and 400v.

Calculate the wavelength associated with the electrons in each case


1
Expert's answer
2020-08-02T15:04:05-0400

a)


"E=\frac{hc}{\lambda}=eV"

"\lambda_1=\frac{hc}{eV_1}=\frac{1239.84}{100}=12.4\ nm"

b)


"\lambda_2=\frac{hc}{eV_2}=\frac{1239.84}{400}=3.1\ nm"